5g of hcl, how much h2 will be produced?
1
Expert's answer
2019-05-06T05:06:27-0400

Zn+2HCl=H2+ZnCl2

n=m/M

M(H2)=2 g/mol

M(HCl)=36.5 g/mol

n (HCl) = 5/36.5=0.14 mol

n (H2) = 1/2 x 0.14=0.07 mol

m (H2)=0.07 x 2 = 0.14 g